# An Extended Closure Relation by LightGBM for Neutrino Radiation Transport in Core-collapse Supernovae

### v1

@@ Abstract
We developed a machine learning model using LightGBM, one of the most popular gradient boosting decision tree methods these days, to predict the Eddington tensor, or the second-order angular moment, for neutrino radiation transport in core-collapse supernova simulations. We use not only zero-th and first moments as in ordinary closure relations but also information on the background matter configuration extensively. For training the model we utilize some post-bounce snapshots from one of our previous Boltzmann radiation-hydrodynamics simulations. The Eddington tensor as well as the zero-th and first angular moments are calculated from the neutrino distribution function obtained in the simulation. LightGBM is light indeed and its high efficiency in training enables us to feed a large number of features and figure out which features are more important than others. We report in this paper the results of the training and validation as well as the generalization of our model: it can reproduce the Eddington factor better in general than the M1-closure relation, one of the most commonly employed algebraic closure relations at present; the generalization performance is also much improved from our previous model based on the deep neural network.
@@ 1 Introduction
Core-collapse supernovae (CCSNe) are explosive demise of massive stars. Despite extensive investigations, the detailed mechanism driving this explosion remains to be established. The neutrino heating (Janka, 2012; Yamada et al., 2024) is the mechanism that most researchers in the society think is most likely at present. It goes as follows. The gravitational collapse of a massive star is halted when the core reaches nuclear density at the center. Then a shock wave is formed and sweeps the accreting outer core, photo-dissociating heavy nuclei therein. The shock energy depleted through this endothermic reaction, leading eventually to the shock stall. In the meantime, a proto-neutron star (PNS) is formed at the center and starts to emit neutrinos copiously. These neutrinos heat up matter behind the shock and re-energize the stagnant shock wave eventually. It should be obvious that the good understanding of neutrino transport is imperative for firmly establishing the theory of the CCSN mechanism.
@@
Numerical simulations have been extensively employed for this purpose (Nagakura et al., 2019a; Mezzacappa et al., 2020; Bollig et al., 2021; Bruenn et al., 2023; Shibagaki et al., 2023; Wang & Burrows, 2024). They comprise the computations of hydrodynamics, gravity, and neutrino transport and reactions. The neutrino transport, the target of the this study, is described by the Boltzmann equation. Its faithful solution, although preferable, is numerically demanding because of the high dimension. The majority of CCSN simulations these days resorts to be established to some approximations. Among them, the truncated moment method is most widely used (Thorne, 1981; Shibata et al., 2011; Cardall et al., 2013; O’Connor, 2015). Instead of the distribution function in momentum space at each spatial point, it considers the angular moments thereof and integrates the Boltzmann equation with respect to the two angles that specify the neutrino momentum, angular moments of the Boltzmann equation, thereby reducing the dimension.
@@
The resultant equations for the moments of different orders are coupled with one another and form an infinite hierarchy. In numerical simulations, it is truncated at some order, typically the first order. The efficacy of this truncated moment method hinges on the accuracy of the closure relation, which one imposes by hand to obtain unsolved moments from the solved ones. If one truncates the hierarchy at the first order, the Eddington tensor, or the second moment of the distribution function, is normally given by the closure relation as a function of the zero-th and first order moments. Among various closure relations proposed in the literature to date, the M1-closure relation is the most popular these days (Levermore, 1984). According to the comparison of the Eddington tensors obtained by the directly integrating the distribution functions obtained in the CCSN simulations that solved the Boltzmann equation faithfully, which we refer to as the Boltzmann simulations hereafter, and the Eddington tensors given by the M1-closure relation for the same data, the diagonal and off-diagonal components of the Eddington tensor may develop discrepancies of several tens of percent and a factor of 2, respectively (Harada et al., 2019; Iwakami et al., 2022).
@@
Some efforts are still being made to enhance the accuracy of closure relation. Nagakura & Johns (2021) suggested a fitting formula for the Eddington factor, the largest eigenvalue of the Eddington tensor, based on the spherically symmetric Boltzmann simulations. Though it offers a reasonably faithful formula, more improvement is required to apply for the multi-dimensional simulations. In a separate endeavor, Harada et al. (2022) developed a tensor basis neural network (TBNN), a deep neural network that provides the Eddington tensor from the neutrino energy density and flux, i.e., the zeroth and first order moments and the local matter velocity. The network was trained by using the data of a Boltzmann simulation. Although the TBNN-closure outperforms the M1-closure in terms of accuracy, there remains room for improvement both in the accuracy and in the generalization performance.
@@
This paper aims at improving the closure relation by machine learning techniques. From our experience in the previous attempt with the TBNN (Harada et al., 2019), we think that a substantial improvement will be achieved only by providing more information other than the two lowest order moments to the machine learning model. In fact, Harada et al. (2019) found that the information on local matter motions was useful as an input to the TBNN. One particularly powerful technique to identify relevant features is the decision tree for regression problems (see Section 2.2 for details). We adopt Light Gradient Boosting Machine (LightGBM) in this paper. It is known that LightGBM can handle a large number of features very efficiently. As one of the decision tree algorithms, on the other hand, it may lack a sufficient smoothness in expressing functions. We attach more importance to the former capability in this study.
@@
This paper is organized as follows. In section 2, we briefly review the M1-closure relation and LightGBM, the machine learning model employed in this paper. Section 3 provides the details of the training: data, feature engineering, validation and hyper parameters. Then, in section 4, we present the main results on the performance of our novel models making comparisons with the Eddington tensors calculated directly from the original data obtained by our Boltzmann simulation as well as with the Eddington tensors derived from the M1-closure relation applied to the zero-th and first angular moments calculated from the same original data. Finally, section 5 is the conclusion. Throughout this paper, we employ the units of $c=1$, where $c$ is the light speed. The metric signature is ($+$)(-$+$)( - + + + ). Greek and Latin indices run over 003333 (spacetime) and 11113333 (spatial).
2 Overviews of the M1-closure relation and LightGBM
@@
The purpose of this paper is to improve by machine learning techniques the estimation of the Eddington tensor in the truncated moment method for the neutrino transport in CCSN simulations. This section is devoted to a brief review of the M1-closure relation in section 2.1 and LightGBM in section 2.2.
2.1 M1-closure relation
@@
The neutrino population in CCSNe is described by the distribution function $f(x,p)$ in phase space, in which $x$ and $p$ are the spacetime coordinate and momentum of neutrinos. It is governed by the Boltzmann equation. The high dimension and stiff source term of this equation poses a numerical challenge under the current computational resources. To the numerical cost, most researchers are opting the truncated moment method nowadays.
@@
where $\varepsilon^{\prime}=-u_{\mu}p^{\prime\mu}$ is the neutrino energy measured in the fluid-rest (FR) frame with $u$ being the fluid four-velocity. By applying appropriate projections to this moment, we can derive the energy densities $E$, energy flux $F^{i}$, and the stress tensors $P^{ij}$ in both the FR and the laboratory (LB) frames:
@@
where $n$ is the unit normal vector to the spatial hypersurface with t = const., $\gamma_{\alpha}{}^{i}=\delta_{\alpha}{}^{i}+n_{\alpha}n^{i}$ is the projection onto that surface with $\delta_{\alpha}{}^{\beta}$ being the Kronecker’s delta, and $h_{\alpha}{}^{i}=\delta_{\alpha}{}^{i}+u_{\alpha}u^{i}$ is the projection onto the plane perpendicular to $u^{\alpha}$. In the truncated moment formalism, the time evolutions of $E_{\rm LB}$ and $F_{\rm LB}^{i}$ are solved normally by setting a closure relation to give $P_{\rm LB}^{ij}$ as a function of $E_{\rm LB}$ and $F_{\rm LB}^{i}$ [^2] In order to solve the spectral evolution of spectrum in general relativity, the third moment of the distribution function is also required. We ignore it in this paper for simplicity, though.. These three quantities are essentially the lowest three angular moments. The exact relations are given, e.g., in Shibata et al. (2011).
@@ 2.2 Light Gradient Boosting Machine
The decision tree (DT) is a machine learning algorithm that performs classification or regression tasks by recursively partitioning a dataset based on certain criteria (Murphy, 2022). Suppose that we find a functional relation between an $n$-dimensional input vector $\boldsymbol{x}$ and an output real number [^3] If the output is integer or discrete number, the task is classification. $y$ from a dataset of size $N$: $\{(\boldsymbol{x}^{i},y^{i})\}_{i=1}^{N}$. In the DT method, each data $(\boldsymbol{x}^{j},y^{j})$ is binary-classified according to $\boldsymbol{x}^{j}$, and the the two destinations are called nodes. The data classified to a certain node is further classified to the daughter nodes recursively. This process hence forms a tree graph. In the DT regression, each node has the prediction value $\hat{y}_{k}=\sum_{j\in D_{k}}y^{j}/|D_{k}|$, where $D_{k}$ is a subset of the data classified to node $k$ and $|D_{k}|$ is its size. The classification criterion at each node is determined so that it would minimize the mean squared error of the prediction value $\hat{y}_{R(L)k}$ for the data in the right (left) daughter node of node $k$. This binary classification is repeated until the error thus obtained gets small enough. Then the output of the DT is the prediction value $\hat{y}^{j}$ of the node, to which the input $\boldsymbol{x}^{j}$ is finally classified.
The Gradient Boosting DT (GBDT, Friedman, 1999) is an improved version of DT. The output of single DT for regression has large errors normally. We can improve the prediction by training another DT for these errors. With recursive correctios with many DTs, the GBDT achieves an accurate prediction. Suppose that $f_{\ell}(\boldsymbol{x})$ is the output of the $\ell$-th DT for the input vector $\boldsymbol{x}$. In the least-square GBDT, each DT is trained so that the following functions
@@
are employed from $\ell=1$; in the above expression, $10<\eta<$ 1 is the learning rate employed to avoid overfitting and $f_{0}(\boldsymbol{x})$ is given conventionally by $\sum_{j=1}^{N}y^{j}/N$. Then, the prediction of GBDT is given by $\hat{y}^{j}=\eta\sum_{\ell=0}^{M}f_{\ell}(\boldsymbol{x}^{j})$, where $M$ is the total number of DTs.
LightGBM (Light Gradient Boosting Machine) is a variant of GBDT introduced by Ke et al. (2017). LightGBM incorporates four techniques to enhance GBDT: (1) leaf-wise tree growth that prioritizes creations of daughter and granddaughter nodes than those of cousin nodes, suppressing the growth of the nodes that focus on irrelervant variables for regression, as shown in figure 1, (2) a histogram-based algorithm, which reduces numerical cost by binning data, (3) the gradient-based one-sided sampling (GOSS) to reduce the training data by focusing on the critical elements, and (4) the exclusive feature bundling, which combines (EFB), which combines features that are not zero simultaneously. Thanks to these techniques, LightGBM excels in computational and memory efficiency compared to traditional models, hence it accelerates the search for the relevant variables of the closure relation.
@@ 3.1 Data Description
The data set used in this study are five snapshots taken from a 2D axisymmetric CCSN simulations performed with the Boltzmann-radiative-hydrodynamics code (Sumiyoshi & Yamada, 2012; Nagakura et al., 2014, 2019a, 2019b). This code solves the Boltzmann equation for neutrinos, the hydrodynamics equations for stellar gas and the Poisson equation for gravity. Simultaneously, we adopt Furusawa–Togashi EOS (Furusawa et al., 2017), which is based on the variational method and is extended to handle a large number of nuclei in the nuclear statistical equilibrium. The progenitor is a non-rotating star with the ZAMS mass of $15\,M_{\odot}$ modeled by Woosley et al. (2002). For the detailed description of this code we refer readers to Sumiyoshi & Yamada (2012), Nagakura et al. (2014), Nagakura et al. (2019a) and Nagakura et al. (2019b). Whereas we consider three neutrino species: electron-type neutrinos and anti-neutrino as well as heavy-lepton-type neutrino, we employ the results for the electron-type neutrino alone in this study. The grid sizes for radius $N_{r}$, zenith angle $N_{\theta}$, energy $N_{\varepsilon}$, momentum angles $N_{\theta_{\nu}}$, $N_{\phi_{\nu}}$ are $(N_{r},N_{\theta},N_{\varepsilon},N_{{\theta}_{\nu}},N_{{\phi}_{\nu}})=(384,128,20,10,6)$. The computational zone covers the region up to 5000 km from the stellar center. In this study, on the other hand, we utilize the data within the ranges of $2000<r\rm{(km)}<$ 200, $0<\theta\rm{(rad)}<\pi$, and $3000<\varepsilon\rm{(MeV)}<$ 300 because this is the region of the greatest relevance for the neutrino heating and shock revival. We extract the distribution function of electron-type neutrino in this region from the simulation data at the post-bounce time of $300\,{\rm ms}$ and use it as the training data. For validation, we employ four other distribution functions sampled at $t=100\,\rm{ms}$, $t=150\,\rm{ms}$, $t=200\,\rm{ms}$ and $t=250\,\rm{ms}$ after bounce. Note that the CCSN simulation we adopt in this paper is identical to what was used in Harada et al. (2022), but the radial extent and the number of snapshots are expanded.
@@
We will construct a machine learning model to give the Eddington tensor as a function of some variables, or features, that include the lower moments. Those we refer to as the basic features are listed in table 1. As discussed later, additional features are generated from these basic ones through feature engineering. The basic and additional features are collectively treated as inputs to the machine.
@@
The output of the machine is the Eddington tensor. In this paper, we do not project it onto the local orthonormal vectors: $e_{r}$, $e_{\theta}$ and $e_{\phi}$, of the spherical coordinates in the LB frame. Instead we employ the orthonormal vectors, $e_{{\rm FP},i}$, one of which is aligned with the neutrino flux and the other two are orthogonal to it and also to each other. We refer to the 3-dimenstional frame defined by these orthonormal vectors as the flux-projected (FP) frame, and the Eddington tensor represented in this frame is denoted as $k^{ij}_{\rm FP}$ hereafter. The FP frame is constructed for each neutrino energy as follows: $e_{{\rm FP},1}$ is the unit vector directed along the neutrino flux. The choice of the other two unit vectors is rather arbitrary. In this paper we obtain $e_{{\rm FP},2}$ from $e_{\theta}$ by subtracting the component parallel to $e_{{\rm FP},1}$, i.e., via the Gram-Schmidt orthogonalization. As a result, $e_{{\rm FP},2}$ so obtained lies in the plane spanned by the flux and $e_{\theta}$. The last vector is derived so that $e_{{\rm FP},1}$, $e_{{\rm FP},2}$ and $e_{{\rm FP},3}$ should form a right-handed orthonormal system. The Eddington tensor in the FP frame is then related with that in the LB frame as $k_{{\rm FP}}^{ij}e_{{\rm FP},i}^{l}e_{{\rm FP},j}^{m}=k_{\rm LB}^{lm}$, where the superscript of $e_{{\rm FR}}$ specifies the components and the repeated indices are summed over.
@@
The reason why we employ the FP frame is the following. The Eddington tensor is not “aligned” in general with the coordinates we choose rather arbitrarily. Since the Eddington tensor is a symmetric second-rank tensor, it becomes diagonal in the appropriately chosen orthonormal frame, which is not equal to the LB frame in general. Suppose that the neutrino distribution in momentum space is axisymmetric with respective to a certain direction misaligned with any one of the orthonormal vectors in the LB frame. Instead the symmetry axis is parallel to the flux vector in this case. In fact, the Eddington tensor is diagonal in the FP frame. It should be stressed that the neutrino distribution in momentum space is not axisymmetric in general, except for the notable case with spherical symmetry. For example, the neutrino flux from a rotating oblate PNS is non-radial and has even nonvanishing $\phi$-component, since neutrinos carry non-zero angular momenta. Then the Eddington tensor in the LB frame has substantial off-diagonal components (Harada et al., 2019; Iwakami et al., 2022). We expect, on the other hand, that the off-diagonal components will be smaller in the FP frame even in this case.
Table 1: Basic Features
Feature Symbol Dimension
Neutrino energy $\varepsilon$ $\rm{(MeV)}$
Optical depth $\tau$
Baryon mass density $\rho_{\text{baryon}}$ $\rm{(g/cm^{3})}$
Energy density (LB) $E_{\text{LB}}$ $\rm{(erg/cm^{3})}$
Flux (LB, $r$)/speed of light $F_{\text{LB}}^{r}$ $\rm{(erg/cm^{3})}$
Flux (LB, $\theta$)/speed of light $F_{\text{LB}}^{\theta}$ $\rm{(erg/cm^{3})}$
Flux (LB, $\phi$)/speed of light $F_{\text{LB}}^{\phi}$ $\rm{(erg/cm^{3})}$
Fluid velocity ($r$) $v^{r}$ $\rm{(cm/s)}$
Fluid velocity ($\theta$) $v^{\theta}$ $\rm{(cm/s)}$
Fluid velocity ($\phi$) $v^{\phi}$ $\rm{(cm/s)}$

### v2

@@ Abstract
We developed a machine learning model using LightGBM, one of the most popular gradient-boosting decision tree methods these days, to predict the Eddington tensor, or the second-order angular moment, for neutrino radiation transport in core-collapse supernova simulations. We use not only the zeroth and first moments of the neutrino distribution function in momentum space as in ordinary closure relations but also information on the background matter configuration extensively. For training the model, we utilize some post-bounce snapshots from one of our previous Boltzmann radiation-hydrodynamics simulations; the Eddington tensor as well as the zeroth and first angular moments are calculated from the neutrino distribution function obtained in the simulation. LightGBM is light indeed, and its high efficiency in training enables us to feed a large number of features and figure out which features are more important than others. In this paper, we report the results of this feature engineering in addition to those of the training, validation, and generalization of our model. We find that the flux factor and non-local features are among the most relevant features; our LightGBM model can reproduce the Eddington factor better in general than the M1 closure relation, one of the most commonly employed algebraic closure relations at present; the generalization performance is also much improved from our previous model based on the deep neural network.
@@ 1 Introduction
Core-collapse supernovae (CCSNe) are explosive demises of massive stars. Despite extensive investigations, the detailed mechanism driving this explosion remains to be established. The neutrino heating (Janka, 2012; Yamada et al., 2024) is the mechanism that most researchers in society believe is most likely at present. The process proceeds as follows: the gravitational collapse of a massive star is halted when the core reaches nuclear density at the center. Then a shock wave is formed and sweeps the accreting outer core, photodissociating heavy nuclei therein. The shock energy is depleted through this endothermic reaction, which eventually leads to a shock stall. In the meantime, a proto-neutron star (PNS) is formed at the center and starts to emit neutrinos copiously. These neutrinos heat up matter behind the shock, re-energize the stagnant shock wave eventually, and revive it. It should be obvious that a good understanding of neutrino transport is imperative to establish the theory of the CCSN mechanism.
@@
Numerical simulations have been extensively employed for this purpose (Nagakura et al., 2019a; Mezzacappa et al., 2020; Bollig et al., 2021; Bruenn et al., 2023; Shibagaki et al., 2023; Wang & Burrows, 2024). They comprise the computations of hydrodynamics, gravity, and neutrino transport and reactions. The neutrino transport, the target of this study, is described by the Boltzmann equation. Its faithful solution, although preferable, is numerically demanding because of its high dimensionality. The majority of CCSN simulations these days resort to some approximations. Among them, the truncated moment method is the most widely used (Thorne, 1981; Shibata et al., 2011; Cardall et al., 2013; O’Connor, 2015). Instead of treating the distribution function in momentum space at each spatial point, it considers the angular moments thereof by integrating the Boltzmann equation with respect to the two angles that specify the neutrino momentum, thereby reducing the dimension.
@@
The resultant equations for the moments of different orders are coupled with one another and form an infinite hierarchy. In numerical simulations, it is truncated at some order, typically the first-order. The efficacy of this truncated moment method hinges on the accuracy of the closure relation, which one imposes by hand to obtain unsolved moments from the solved ones. If one truncates the hierarchy at the first-order, the Eddington tensor, or the second moment of the distribution function, is normally given by the closure relation as a function of the zeroth and first-order moments. Among various closure relations proposed in the literature to date, the M1 closure relation is the most popular these days (Levermore, 1984). According to the comparison of the Eddington tensors derived by the direct integration of the distribution functions obtained in the CCSN simulations that solve the Boltzmann equation faithfully, which we refer to as the Boltzmann simulations hereafter, and the Eddington tensors given by the M1 closure relation for the same data, the diagonal and off-diagonal components of the Eddington tensor may develop discrepancies of several tens of percent and a factor of 2, respectively (Harada et al., 2019; Iwakami et al., 2022).
@@
Some efforts are still being made to enhance the accuracy of closure relations. Nagakura & Johns (2021) suggested a fitting formula for the Eddington factor, the largest eigenvalue of the Eddington tensor, based on the spherically symmetric Boltzmann simulations. Though it offers a fairly good formula, more improvement is desirable to apply it to the multi-dimensional simulations. In a separate endeavor, Harada et al. (2022) developed a tensor basis neural network (TBNN), a deep neural network that provides the Eddington tensor from the neutrino energy density and flux, i.e., the zeroth and first-order moments and the local matter velocity. The network was trained by using the data of a Boltzmann simulation. Although the TBNN-closure outperforms the M1 closure in terms of accuracy, there remains room for improvement both in the accuracy and in the generalization performance.
@@
This paper aims to improve the closure relation by machine learning techniques. From our experience in the previous attempt with the TBNN (Harada et al., 2019), we think that a substantial improvement will be achieved only by providing more information other than the two lowest order moments to the machine learning model. In fact, Harada et al. (2019) found that the information on local matter motions was useful as an input to the TBNN. One particularly powerful technique to identify relevant features is the decision tree for regression problems (see Section 2.2 for details). We adopt Light Gradient Boosting Machine (LightGBM) in this paper. It is known that LightGBM can handle a large number of features very efficiently. As one of the decision tree algorithms, on the other hand, it may lack sufficient smoothness in expressing functions. We attach more importance to the former capability in this study.
@@
This paper is organized as follows. In section 2, we briefly review the M1 closure relation and LightGBM, the machine learning model employed in this paper. Section 3 provides the details of the training: data, feature engineering, validation, and hyperparameters. Then, in section 4, we present the main results on the performance of our novel models making comparisons with the Eddington tensors calculated directly from the original data obtained by our Boltzmann simulation as well as with the Eddington tensors derived from the M1 closure relation applied to the zeroth and first angular moments calculated from the same original data. Finally, section 5 is the conclusion. Throughout this paper, we employ the units of $c=1$, where $c$ is the light speed. The metric signature is ($+$)(-$+$)( - + + + ). Greek and Latin indices run over 003333 (spacetime) and 11113333 (spatial).
2 Overviews of the M1 closure relation and LightGBM
@@
The purpose of this paper is to improve by machine learning techniques the estimation of the Eddington tensor in the truncated moment method for the neutrino transport in CCSN simulations. This section is devoted to a brief review of the M1 closure relation in section 2.1 and LightGBM in section 2.2.
2.1 M1 closure relation
@@
The neutrino population in CCSNe is described by the distribution function $f(x,p)$ in phase space, in which $x$ and $p$ are the spacetime coordinates and momentum of neutrinos. It is governed by the Boltzmann equation. The high dimensionality and stiff source terms of this equation pose numerical challenges with current computational resources. Because of its high numerical cost, most researchers are opting for the truncated moment method nowadays.
@@
where $\varepsilon^{\prime}=-u_{\mu}p^{\prime\mu}$ is the neutrino energy measured in the fluid-rest (FR) frame with $u$ being the fluid four-velocity. By applying appropriate projections to this moment, we can derive the energy densities $E$, energy flux $F^{i}$, and the stress tensors $P^{ij}$ in both the FR and laboratory (LB) frames:
@@
where $n$ is the unit normal vector to the spatial hypersurface with $t=\rm{const.}$; $\gamma_{\alpha}{}^{i}=\delta_{\alpha}{}^{i}+n_{\alpha}n^{i}$ is the projector onto that surface with $\delta_{\alpha}{}^{\beta}$ being the Kronecker’s delta, whereas $h_{\alpha}{}^{i}=\delta_{\alpha}{}^{i}+u_{\alpha}u^{i}$ is the projector onto the plane perpendicular to $u^{\alpha}$. In the truncated moment formalism, the time evolutions of $E_{\rm LB}$ and $F_{\rm LB}^{i}$ are solved normally by setting a closure relation to give $P_{\rm LB}^{ij}$ as a function of $E_{\rm LB}$ and $F_{\rm LB}^{i}$ [^2] In order to solve the evolution of spectrum in general relativity, the third moment of the distribution function is also required. We ignore it in this paper for simplicity, though.. These three quantities are essentially the lowest three angular moments. Their exact relations are given, e.g., in Shibata et al. (2011).
@@ 2.2 Light Gradient Boosting Machine
The decision tree (DT) is a machine learning algorithm that performs classification or regression tasks by recursively partitioning a dataset based on certain criteria (Murphy, 2022). Suppose that we find a functional relation between an $n$-dimensional input vector $\boldsymbol{x}$ and an output real number [^3] If the output is integer or discrete number, the task is a classification. $y$ from a dataset of size $N$: $\{(\boldsymbol{x}^{i},y^{i})\}_{i=1}^{N}$. In the DT method, each data $(\boldsymbol{x}^{j},y^{j})$ is binary-classified according to $\boldsymbol{x}^{j}$, and the two destinations are called nodes. The data classified to a certain node is further classified to one of its two daughter nodes recursively. This process hence forms a tree graph. In the DT regression, each node has the prediction value $\hat{y}_{k}=\sum_{j\in D_{k}}y^{j}/|D_{k}|$, where $D_{k}$ is a subset of the data classified to node $k$ and $|D_{k}|$ is its size. The classification criterion at each node is determined so that it would minimize the mean squared error of the predicted value $\hat{y}_{R(L)k}$ at the right (left) daughter node of node $k$. This binary classification is repeated until the error thus obtained gets small enough. Then the output of the DT is the prediction value $\hat{y}^{j}$ of the node, to which the input $\boldsymbol{x}^{j}$ is finally classified.
The Gradient Boosting DT (GBDT, Friedman, 1999) is an improved version of DT. The output of a single DT for regression has large errors normally. We can improve a prediction by training another DT for these errors. With recursive corrections with many DTs, the GBDT achieves an accurate prediction. Suppose that $f_{\ell}(\boldsymbol{x})$ is the output of the $\ell$-th DT for the input vector $\boldsymbol{x}$. In the least-square GBDT, each DT is trained so that the following functions
@@
should be minimized from $\ell=1$. In the above expression, $10<\eta<$ 1 is the learning rate employed to avoid overfitting and $f_{0}(\boldsymbol{x})$ is given conventionally by $\sum_{j=1}^{N}y^{j}/N$. Then, the prediction of GBDT is given by $\hat{y}^{j}=\eta\sum_{\ell=0}^{M}f_{\ell}(\boldsymbol{x}^{j})$, where $M$ is the total number of DTs.
LightGBM (Light Gradient Boosting Machine) is a variant of GBDT introduced by Ke et al. (2017). LightGBM incorporates four techniques to enhance GBDT: (1) leaf-wise tree growth that prioritizes creations of daughter and granddaughter nodes than those of cousin nodes, suppressing the growth of the nodes that focus on irrelevant variables for regression, as shown in figure 1, (2) a histogram-based algorithm, which reduces numerical cost by binning data, (3) the gradient-based one-sided sampling (GOSS) to reduce the training data by focusing on the critical elements, and (4) the exclusive feature bundling (EFB), which combines features that are not zero simultaneously. Thanks to these techniques, LightGBM excels in computational and memory efficiency in the training, compared to traditional models. It can hence accelerate the search for relevant variables in the closure relation.
@@ 3.1 Data Description
The data set used in this study is composed of five snapshots taken from a 2D axisymmetric CCSN simulation performed with the Boltzmann-radiation-hydrodynamics code (Sumiyoshi & Yamada, 2012; Nagakura et al., 2014, 2019a, 2019b). This code solves the Boltzmann equation for neutrinos, the hydrodynamics equations for stellar gas, and the Poisson equation for gravity simultaneously. We adopt Furusawa–Togashi EOS (Furusawa et al., 2017), which is based on the variational method and is extended to handle a large number of nuclei in the nuclear statistical equilibrium. The progenitor is a non-rotating star with the ZAMS mass of $15\,M_{\odot}$ modeled by Woosley et al. (2002). For a detailed description of this code we refer readers to Sumiyoshi & Yamada (2012), Nagakura et al. (2014), Nagakura et al. (2019a), and Nagakura et al. (2019b). Whereas we consider three neutrino species: electron-type neutrinos and anti-neutrinos as well as heavy-lepton-type neutrinos, we employ the results for the electron-type neutrino alone in this study. The grid sizes for radius $N_{r}$, zenith angle $N_{\theta}$, energy $N_{\varepsilon}$, momentum angles $N_{\theta_{\nu}}$, $N_{\phi_{\nu}}$ are $(N_{r},N_{\theta},N_{\varepsilon},N_{{\theta}_{\nu}},N_{{\phi}_{\nu}})=(384,128,20,10,6)$. The computational zone covers the region up to 5000 km from the stellar center, with $0<\theta\rm{(rad)}<\pi$ and $3000<\varepsilon\rm{(MeV)}<$ 300. In this study, on the other hand, we utilize the data within the ranges of $2000<r\rm{(km)}<$ 200 because this is the region of the greatest relevance for the neutrino heating and shock revival. We extract the distribution function of electron-type neutrino in this region from the simulation data at the post-bounce time of $300\,{\rm ms}$ and use it as the training data. For validation, we employ four other distribution functions sampled at $t=100\,\rm{ms}$, $t=150\,\rm{ms}$, $t=200\,\rm{ms}$ and $t=250\,\rm{ms}$ after bounce. Note that the CCSN simulation we adopt in this paper is identical to what was used in Harada et al. (2022), but the radial extent and the number of snapshots are expanded.
@@
We will construct a machine learning model to give the Eddington tensor as a function of some variables, or features, that include the lower moments. Those we refer to as the basic features includes the variables to construct the conventional closure relations. As discussed later, additional features are generated from these basic ones through feature engineering. The basic and additional features are collectively treated as inputs to the machine.
@@
The output of the machine is the Eddington tensor. In this paper, we do not project it onto the local orthonormal vectors: $e_{r}$, $e_{\theta}$, and $e_{\phi}$, of the spherical coordinates in the LB frame. Instead, we employ the orthonormal vectors, $e_{{\rm FP},i}$, one of which is aligned with the neutrino flux, and the other two are orthogonal to it and also to each other. We refer to the 3-dimensional frame defined by these orthonormal vectors as the flux-projected (FP) frame, and the Eddington tensor represented in this frame is denoted as $k^{ij}_{\rm FP}$ hereafter. The FP frame is constructed for each neutrino energy as follows: $e_{{\rm FP},1}$ is the unit vector parallel to the neutrino flux. The choice of the other two unit vectors is rather arbitrary. In this paper we obtain $e_{{\rm FP},2}$ from $e_{\theta}$ by subtracting the component parallel to $e_{{\rm FP},1}$, i.e., via the Gram-Schmidt orthogonalization. As a result, $e_{{\rm FP},2}$ so obtained lies in the plane spanned by the flux and $e_{\theta}$. The last vector is derived so that $e_{{\rm FP},1}$, $e_{{\rm FP},2}$ and $e_{{\rm FP},3}$ should form a right-handed orthonormal system. The Eddington tensor in the FP frame is then related to that in the LB frame as $k_{{\rm FP}}^{ij}e_{{\rm FP},i}^{l}e_{{\rm FP},j}^{m}=k_{\rm LB}^{lm}$, where the superscript of $e_{{\rm FP}}$ specifies the components and the repeated indices are summed over.
@@
The reason why we employ the FP frame is the following. The Eddington tensor is not “aligned” in general with the coordinates we choose rather arbitrarily. Since the Eddington tensor is a symmetric second-rank tensor, it becomes diagonal in the appropriately chosen orthonormal frame, which is not equal to the LB frame in general. Suppose that the neutrino distribution in momentum space is axisymmetric with respect to a certain direction misaligned with any one of the orthonormal vectors in the LB frame. Instead, the symmetry axis is parallel to the flux vector in this case. In fact, the Eddington tensor is diagonal in the FP frame. It should be stressed that the neutrino distribution in momentum space is not axisymmetric in general, except for the notable case of spherical symmetry. For example, the neutrino flux from a rotating oblate PNS in space is non-radial and has even nonvanishing $\phi$-component, since neutrinos carry non-zero angular momenta. Then the Eddington tensor in the LB frame has substantial off-diagonal components (Harada et al., 2019; Iwakami et al., 2022). We expect, on the other hand, that the off-diagonal components will be smaller in the FP frame even in this case.